# An iterative method for the solution of Laplace-like equations in high and very high space dimensions

### v3

@@ 3 A moment-based analysis
The singular values of the matrices $T$ and $T^{t}$, respectively, can in general only be determined numerically. The matrices $T$ assigned to complete graphs, in which all vertices are connected with each other, are an exception. It is not difficult to see that the associated matrices $S=T^{t}T$ have the eigenvalue one of multiplicity one and the eigenvalue $(m+2)/2$ of multiplicity $m-1$ in terms of the number $m$ of vertices. The square roots of these eigenvalues are the singular values of $T$.
Orthogonal projections, matrices $T^{t}=P$ of dimension $(m\times n)$ with one as the only singular value, are another important exception. The quantities (8) take the value $m$ in this case. The expected value and the variance are therefore
$E=\frac{m}{n},\quad V=\frac{m}{n}\left(1-\frac{m}{n}\right)\frac{2}{n+2},$ (25)
or after rescaling, that is, after multiplying $P$ by the factor $\sqrt{n/m}$,
$E=1,\quad V=\frac{n-m}{n+2}\,\frac{2}{m}.$ (26)
Since $\|\Sigma_{0}x\|^{2}=\|x\|^{2}$ for orthogonal projections, for these matrices the quotient (24) follows the beta distribution with the probability density function
$f(t)=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}\;t^{\alpha-1}(1-t)^{\beta-1}$ (27)
with parameters $\alpha=m/2$ and $\beta=(n-m)/2$; see (Abramowitz-Stegun, , Eq. 26.5.3). In a context similar to ours, this has already been observed in Frankl-Maehara . Orthogonal projections are accessible to a complete analysis. The next section is therefore devoted to their study.
4 Orthogonal projections and measure concentration
We begin with a simple, but very helpful corollary from Lemma 4.
Lemma 8
Let $A$ be an arbitrary matrix of dimension $m\times n$, $m<n$, with singular value decomposition $A=Q\Sigma\widetilde{Q}^{t}$. Then the probabilities
($\mathbb{P}\big{(}\big{\{}x\,\big{|}\,\|Ax\|^{2}\leq\delta\|A\|^{2}\|x\|^{2}\big{\}}\big{)},\quad\mathbb{P}\big{(}\big{\{}x\,\big{|}\,\|\Sigma x\|^{2}\leq\delta\|\Sigma\|^{2}\|x\|^{2}\big{\}}\big{)}$ (1)
are equal and the first one depends only on the singular values of $A$.
Proof
This follows from the invariance of these probabilities under orthogonal transformations. Let $\chi(x)=1$ if $\|Ax\|^{2}\leq\delta\|A\|^{2}\|x\|^{2}$ holds and $\chi(x)=0$ otherwise. The transformed function, with the values $\chi(\widetilde{Q}x)$, is then the characteristic function of the set of all $x$ for which $\|\Sigma x\|^{2}\leq\delta\|\Sigma\|^{2}\|x\|^{2}$ holds. ∎
The next theorem is an only slightly modified version of Theorem 2.4 in Yserentant_2022 . Its proof is based on the observation that the probability (1) is equal to the volume of the set of all $x$ inside the unit ball for which $\|Ax\|^{2}\leq\delta\|A\|^{2}\|x\|^{2}$ holds, up to the division by the volume of the unit ball itself. This is an implication of Lemma 3 if the characteristic function of the unit ball is chosen as the function $W$.
Theorem 4.1
Let the $(m\times n)$-matrix $P$ be an orthogonal projection. Then
($\mathbb{P}\big{(}\big{\{}x\,\big{|}\,\|Px\|^{2}\leq\delta\|x\|^{2}\big{\}}\big{)}=F(\delta),\quad 0\leq\delta<1,$ (2)
holds, where the function $F(\delta)=F(m,n;\delta)$ is defined by the integral expression
$F(\delta)=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}\,\int_{0}^{\delta}t^{\alpha-1}(1-t)^{\beta-1}\,\mathrm{d}t$ (3)
and the two exponents $\alpha$ and $\beta$ are given by
$\alpha=\frac{m}{2},\quad\beta=\frac{n-m}{2}.$ (4)
The function (2) tends to the limit value $F(1)=1$ as $\delta$ goes to one.
Proof
By Lemma 8, we can restrict ourselves to the matrix $P$ that extracts from a vector in $\mathbb{R}^{n}$ its first $m$ components. We split the vectors in $\mathbb{R}^{n}$ into a part $x\in\mathbb{R}^{m}$ and a remaining part $y\in\mathbb{R}^{n-m}$. The set whose volume has to be calculated according to the above remark, then consists of the points in the unit ball for which
$\|x\|^{2}\leq\delta\big{(}\,\|x\|^{2}+\|y\|^{2}\,\big{)}$
or, resolved for the norm of the component $x\in\mathbb{R}^{m}$,
$\|x\|\leq\varepsilon\,\|y\|,\quad\varepsilon=\sqrt{\frac{\delta}{1-\delta}}\,,$
holds. This volume can be expressed as double integral
$\int\left(\int H\big{(}\varepsilon\|y\|-\|x\|\big{)}\chi\big{(}\|x\|^{2}+\|y\|^{2}\big{)}\,\mathrm{d}x\right)\!\,\mathrm{d}y,$
where $H(t)=0$ for $t<0$, $H(t)=1$ for $t\geq 0$, $\chi(t)=1$ for $t\leq 1$, and $\chi(t)=0$ for arguments $t>1$. If $\delta$ tends to one and $\varepsilon$ thus to infinity, the value of this double integral tends, by the dominated convergence theorem, to the volume of the unit ball and the expression (2) tends to one, as claimed. In terms of polar coordinates, that is, according to (14), the double integral is given by
$(n-m)\nu_{n-m}\int_{0}^{\infty}\left(m\nu_{m}\int_{0}^{\varepsilon s}\chi\big{(}r^{2}+s^{2}\big{)}r^{m-1}\,\mathrm{d}r\right)s^{n-m-1}\,\mathrm{d}s,$
where $\nu_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$. By substituting $t=r/s$ in the inner integral, the upper bound becomes independent of $s$ and the integral simplifies to
$(n-m)\nu_{n-m}\int_{0}^{\infty}\left(m\nu_{m}\,s^{m}\!\int_{0}^{\varepsilon}\chi\big{(}s^{2}(1+t^{2})\big{)}t^{m-1}\,\mathrm{d}t\right)s^{n-m-1}\,\mathrm{d}s.$
Interchanging the order of integration, it attains the value
$\frac{(n-m)\nu_{n-m}\,m\nu_{m}}{n}\int_{0}^{\varepsilon}\frac{t^{m-1}}{(1+t^{2})^{n/2}}\,\mathrm{d}t.$
With $\varphi(t)=t^{2}/(1+t^{2})$ and the constants $\alpha$ and $\beta$ from (4), one obtains
$\int_{0}^{\varepsilon}\frac{t^{m-1}}{(1+t^{2})^{n/2}}\,\mathrm{d}t=\frac{1}{2}\int_{0}^{\varepsilon}\varphi(t)^{\alpha-1}(1-\varphi(t))^{\beta-1}\varphi^{\prime}(t)\,\mathrm{d}t$
and, because of $\varphi(0)=0$ and $\varphi(\varepsilon)=\delta$, therefore the representation
$\int_{0}^{\varepsilon}\frac{t^{m-1}}{(1+t^{2})^{n/2}}\,\mathrm{d}t=\frac{1}{2}\int_{0}^{\delta}t^{\alpha-1}(1-t)^{\beta-1}\,\mathrm{d}t$
of the integral. Dividing the expression above by $\nu_{n}$ and recalling that
$\nu_{d}=\frac{2}{d}\,\frac{\pi^{d/2}}{\Gamma(d/2)},$
this completes the proof of the theorem. ∎
Figure 2: The frequency distribution of the values $\|Px\|^{2}$ for the dimensions $m=30$ and $n=165$ together with the exact distribution from Theorem 4.1 and its approximation by the Gauss function (23)
That is, for points $x$ on the unit sphere, the values $\|Px\|^{2}$ follow the beta distribution mentioned at the end of the last section. For smaller dimensions, the frequency distribution $f(t)=F^{\prime}(t)$ and its approximation by the Gauss function (23) still differ considerably, as can be seen in Fig. 2. This difference vanishes with increasing dimensions. In fact, it is common knowledge in statistics that beta distributions behave in the limit of large parameters almost like normal distributions. If a random variable $X(\omega)$ with expected value $E$ and standard deviation $\sigma$ follows a beta distribution with parameters $\alpha$ and $\beta$, the distribution of the random variable
$\frac{X(\omega)-E}{\sigma}$ (5)
differs the less from the standard normal distribution as $\alpha$ and $\beta$ become larger. For the sake of completeness, we prove this in the following. Let
$B(\alpha,\beta;x)=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}\int_{0}^{x}t^{\alpha-1}(1-t)^{\beta-1}\,\mathrm{d}t$ (6)
be the distribution function of the beta distribution with parameters $\alpha$ and $\beta$, let $E$ and $V$ be the assigned expected value and the assigned variance given by
$E=\frac{\alpha}{\alpha+\beta},\quad V=\frac{\alpha\beta}{(\alpha+\beta)^{2}(\alpha+\beta+1)},$ (7)
and let $\sigma=\sqrt{V}$ be the corresponding standard deviation. The function (6) is first only defined for arguments $0\leq x\leq 1$. To avoid problems with its domain of definition, we extend it by the value zero for arguments $x<0$ and the value one for arguments $x>1$. If $\alpha$ and $\beta$ are greater than one, which is the only us interesting case, the extended function is a continuously differentiable cumulative distribution function with a density that vanishes outside the interval $10<t<$ 1. Let
$\widetilde{B}(\alpha,\beta;x)=B(\alpha,\beta;\sigma x+E)$ (8)
be the accordingly rescaled distribution function.
Theorem 4.2
The function (8) tends pointwise to the distribution function
$\Phi(x)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}\mathrm{e}^{-t^{2}/2}\,\mathrm{d}t.$ (9)
of the standard normal distribution as $\alpha$ and $\beta$ go to infinity.
Proof
Before we start with the proof itself, we recall that it suffices to show that the assigned densities converge pointwise to the density of the normal distribution. This may be surprising in view of the monotone or the dominated convergence theorem, but is easily explained and is often utilized in stochastics. Let $f_{1},f_{2},\ldots$ be a sequence of densities that converges pointwise to a density $f$. The nonnegative functions
$h_{n}(x)=\inf_{k\geq n}f_{k}(x)$
then tend pointwise monotonously from below to $f$ and it is
[…]
where we have used that the integral of $f_{n}$ over $\mathbb{R}$ takes the value one. The two outer integrals converge, by the monotone convergence theorem, to the limits
[…]
This proves that the middle integral converges to the limit
$\lim_{n\to\infty}\,\int_{-\infty}^{x}\!f_{n}(t)\,\mathrm{d}t\,=\int_{-\infty}^{x}\!f(t)\,\mathrm{d}t.$
Pointwise convergence of the densities thus implies convergence in distribution.
Now let us fix an arbitrary finite interval and assume that the parameters $\alpha$ and $\beta$ are already so large that, for the points $x$ in this interval, the quantities
$\bigg{(}\frac{\sigma}{E}\,x\bigg{)}^{2}\!=\,\frac{1}{\alpha}\,\frac{\beta x^{2}}{\alpha+\beta+1},\quad\bigg{(}\frac{\sigma}{1-E}\,x\bigg{)}^{2}\!=\,\frac{1}{\beta}\,\frac{\alpha x^{2}}{\alpha+\beta+1}$
take values less than $9/100$. On the given interval, the density of the distribution (8), its derivative with respect to the variable $x$, can then be written as follows
$\frac{\Gamma(\alpha+\beta)(E+\sigma x)^{\alpha-1}(1-E-\sigma x)^{\beta-1}\sigma}{\Gamma(\alpha)\Gamma(\beta)}.$
To proceed, we need Stirling’s formula (see (DLMF, , Eq. 5.6.1)), the representation
$\Gamma(z)=\sqrt{2\pi}\,z^{z-1/2}\mathrm{e}^{-z}\,\mathrm{e}^{\mu(z)},\quad 0<\mu(z)<\frac{1}{12z},$
of the gamma function for real arguments $z>0$. Introducing the function
$g(z)=\ln(\sqrt{2\pi})+\bigg{(}z-\frac{1}{2}\bigg{)}\ln(z)-z,$
the logarithm of the density breaks down into the sum of the three terms
$A=\ln(\sigma)+g(\alpha+\beta)-g(\alpha)-g(\beta)+(\alpha-1)\ln(E)+(\beta-1)\ln(1-E),$
$B=(\alpha-1)\ln\bigg{(}1+\frac{\sigma}{E}\,x\bigg{)}+(\beta-1)\ln\bigg{(}1-\frac{\sigma}{1-E}\,x\bigg{)},$
$C=\mu(\alpha+\beta)-\mu(\alpha)-\mu(\beta).$
The third term $C$ tends by definition to zero as $\alpha$ and $\beta$ go to infinity. The first of the three terms can be easily calculated and reduces to the value
$A=\ln\bigg{(}\frac{1}{\sqrt{2\pi}}\bigg{)}+\frac{1}{2}\ln\bigg{(}\frac{\alpha+\beta}{\alpha+\beta+1}\bigg{)}.$
By means of the second-order Taylor expansion
$\ln(1+t)=P(t)+R(t),\quad P(t)=t-\frac{1}{2}\,t^{2},$
of $\ln(1+t)$ around $t=0$, the second term splits into the polynomial part
$B_{1}=(\alpha-1)P\bigg{(}\frac{\sigma}{E}\,x\bigg{)}+(\beta-1)P\bigg{(}\!-\frac{\sigma}{1-E}\,x\bigg{)}$
and a remainder $B_{2}$. The polynomial part possesses the representation
$B_{1}={}-\frac{1-a}{2}\,x^{2}+bx,$
with coefficients $a$ and $b$ that depend on the parameters $\alpha$ and $\beta$ and are given by
$a\,=\frac{1}{\alpha}\,\frac{\beta}{\alpha+\beta+1}+\frac{1}{\alpha+\beta+1}+\frac{1}{\beta}\,\frac{\alpha}{\alpha+\beta+1},$
$\;b\,=\bigg{(}\frac{1}{\beta}\,\frac{\alpha}{\alpha+\beta+1}\bigg{)}^{1/2}\!\!-\;\bigg{(}\frac{1}{\alpha}\,\frac{\beta}{\alpha+\beta+1}\bigg{)}^{1/2}.$
Because of $|R(t)|\leq|t|^{3}$ for $|t|\leq 3/10$ and our assumption about the size of $\alpha$ and $\beta$ in dependence of the interval under consideration, the term $B_{2}$ can be estimated as
$|B_{2}|\leq(\alpha-1)\bigg{(}\frac{x^{2}}{\alpha}\bigg{)}^{3/2}\!\!+\;(\beta-1)\bigg{(}\frac{x^{2}}{\beta}\bigg{)}^{3/2}.$
The logarithm of the density tends on the given interval therefore uniformly to
$\ln\bigg{(}\frac{1}{\sqrt{2\pi}}\bigg{)}-\,\frac{1}{2}\,x^{2}$
as $\alpha$ and $\beta$ go to infinity. This implies locally uniform convergence of the densities themselves and thus at least pointwise convergence of the distributions ∎
However, the key message of this section remains that the values $\|Px\|^{2}$ cluster, for points $x$ on the unit sphere, more and more around their expected value $\xi=m/n$ as the dimensions increase. Much better than by the previous theorem this is reflected by the following result, which again is more or less taken from Yserentant_2022 . A similar technique, based in the same way on the Markov inequality, was used in Dasgupta-Gupta . The following estimates are expressed in terms of the function
$\phi(\vartheta)=\sqrt{\vartheta}\,\exp\left(\frac{1-\vartheta}{2}\right).$ (10)
It increases strictly on the interval $0\leq\vartheta\leq 1$, reaches its maximum value one at the point $\vartheta=1$, and decreases strictly from there.
Theorem 4.3
Let the $(m\times n)$-matrix $P$ be an orthogonal projection and let $\xi$ be ratio of the two dimensions $m$ and $n$. For $0<\delta<\xi$, then
($\mathbb{P}\big{(}\big{\{}x\,\big{|}\,\|Px\|^{2}\leq\delta\,\|x\|^{2}\big{\}}\big{)}\leq\,\phi\left(\frac{\delta}{\xi}\right)^{m}\phantom{.}$ (11)
holds. For $\xi<\delta<1$, this estimate is complemented by the estimate
($\mathbb{P}\big{(}\big{\{}x\,\big{|}\,\|Px\|^{2}>\delta\,\|x\|^{2}\big{\}}\big{)}\leq\,\phi\left(\frac{\delta}{\xi}\right)^{m}.$ (12)
Proof
As before, we can restrict ourselves to the matrix $P$ that extracts from a vector in $\mathbb{R}^{n}$ its first $m$ components. The characteristic function $\chi$ of the set of all $x$ for which the estimate $\|Px\|^{2}\leq\delta\,\|x\|^{2}$ holds satisfies, for all $t>0$, the estimate
$\chi(x)\leq\exp\big{(}\,t\,\big{(}\delta\|x\|^{2}-\|Px\|^{2}\big{)}\big{)}$
by a product of univariate functions. Choosing the normed weight function (5), the left-hand side of (11) can, by Lemma 3, therefore be estimated by the integral
$\left(\frac{1}{\sqrt{\pi}}\right)^{n}\!\int\exp\big{(}\,t\,\big{(}\delta\|x\|^{2}-\|Px\|^{2}\big{)}\big{)}\exp\big{(}-\|x\|^{2}\big{)}\,\mathrm{d}x,$
which remains finite for all $t$ in the interval $0\leq t<1/\delta$. It splits into a product of one-dimensional integrals and takes, for given $t$, the value
$\left(\frac{1}{1-\delta t+t}\right)^{m/2}\left(\frac{1}{1-\delta t}\right)^{(n-m)/2}.$
This expression attains its minimum on the interval $0<t<1/\delta$ at
$t=\frac{\xi-\delta}{(1-\delta)\delta}.$
At this point $t$ it takes the value
$\left(\left(\frac{\delta}{\xi}\right)^{1/2}\!\left(\frac{1-\delta}{1-\xi}\right)^{\gamma}\;\right)^{m},\quad\gamma=\frac{1-\xi}{2\xi}.$
With the abbreviation $\delta/\xi=\vartheta$ and because of $\vartheta\xi<1$ and $\xi<1$, the logarithm
$\ln\left(\left(\frac{1-\delta}{1-\xi}\right)^{\gamma}\;\right)=\,\frac{1-\xi}{\xi}\,\ln\left(\frac{1-\vartheta\xi}{1-\xi}\right)$
possesses the power series expansion
$\frac{1-\vartheta}{2}-\frac{1}{2}\,\sum_{k=1}^{\infty}\left(\frac{1-\vartheta^{k}}{k}-\,\frac{1-\vartheta^{k+1}}{k+1}\right)\xi^{k}.$
Because the series coefficients are for all $\vartheta\geq 0$ greater than or equal to zero and, by the way, polynomial multiples of $(1-\vartheta)^{2}$, the estimate (11) follows.
The proof of the second estimate is almost identical to the proof of the first. For sufficiently small $t>0$, the expression (12) can be estimated by the integral
$\left(\frac{1}{\sqrt{\pi}}\right)^{n}\!\int\exp\big{(}\,t\,\big{(}\|Px\|^{2}-\delta\|x\|^{2}\big{)}\big{)}\exp\big{(}-\|x\|^{2}\big{)}\,\mathrm{d}x.$
This integral splits into a product of one-dimensional integrals and takes the value
$\left(\frac{1}{1+\delta t-t}\right)^{m/2}\left(\frac{1}{1+\delta t}\right)^{(n-m)/2},$
which attains, for $\delta<1$, on the interval $0<t<1/(1-\delta)$ its minimum at
$t=\frac{\delta-\xi}{(1-\delta)\delta}.$
At this point $t$ it again takes the value
$\left(\left(\frac{\delta}{\xi}\right)^{1/2}\!\left(\frac{1-\delta}{1-\xi}\right)^{\gamma}\;\right)^{m},\quad\gamma=\frac{1-\xi}{2\xi}.$
This leads, as above, to the estimate (12). ∎
If the dimension ratio $\xi=m/n$ is fixed or only tends to a limit value $\delta_{0}$, then the probability distributions (2) and the functions (3), respectively, tend to a step function with jump discontinuity at $\delta_{0}$. Figure 3 reflects this behavior. One can look at the estimates from Theorem 4.3 also from a different perspective and consider the rescaled counterparts $P^{\prime}=\sqrt{n/m}\,P$ of the projections $P$. The distributions
($\mathbb{P}\big{(}\big{\{}x\,\big{|}\,\|P^{\prime}x\|^{2}\leq\delta\|x\|^{2}\big{\}}\big{)}$ (13)
assigned to them tend, outside every small neighborhood of $\delta_{0}=1$, exponentially to a step function with jump discontinuity at $\delta_{0}=1$ as $m$ goes to infinity, independent of the ratio $m/n$ of the two dimensions, no matter how small it may be.
Figure 3: The probability distributions (2) as function of $10\leq\delta<$ 1 for $m=2^{k}$, $k=1,\ldots,16$, and $n=2m$
@@ 5 The iterative procedure
and the functions $P_{k}(\lambda)$ are polynomials of order $k$ with value $P_{k}(0)=1$. Throughout this section, we assume that the expression $\|T^{t}\eta\|^{2}$ possesses the expected value one. We restrict ourselves to the analysis of this iteration in the spectral Barron spaces $\mathcal{B}_{s}$ equipped with the norm (16), or (15) in radial-angular representation. In a corresponding sense, this analysis can be transferred to the Hilbert spaces $H^{s}$.
@@
If the solution $U$ lies in the Barron space $\mathcal{B}_{s}(\mathbb{R}^{n})$, $s\geq 0$, this also holds for the iterates $U_{k}$. For all coefficients $\mu\geq 0$, the norm (16) of the error, given by
$\|U-U_{k}\|_{s}=\int\big{|}P_{k}\big{(}\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)\big{)}\widehat{U}(\omega)\big{|}\|\omega\|^{s}\,\mathrm{d}\omega,$ (4)
@@
be the to the interval $a\leq\lambda\leq b$ transformed Chebyshev polynomial $T_{k}$ of the first kind of degree $k$. The norm (16) of the iteration error (2) then satisfies the estimate
$\|U_{k}-U\|_{s}\leq\frac{2r^{k}}{1+r^{2k}}\,\|U\|_{s}+\|U-U_{\delta}\|_{s},\quad r=\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1},$ (6)
@@
$\delta\leq\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)$ (7)
@@
This follows from $0\leq\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)\leq b$ for $\omega\in\mathbb{R}^{n}$ and the estimates
$|P_{k}(\lambda)|\leq 1\text{\;for $0\leq\lambda\leq a$},\quad|P_{k}(\lambda)|\leq\frac{2r^{k}}{1+r^{2k}}\text{\;for $a\leq\lambda\leq b$}$
@@
Depending on the size of $\kappa=\|T^{t}\|^{2}/\delta$, the norm of the error $U-U_{k}$ soon reaches the size of the norm of $U-U_{\delta}$. The idea behind the estimate (6) is that in high space dimensions the condition (7) is satisfied for nearly all $\omega$ and that the part $U-U_{\delta}$ of the solution $U$ is therefore often negligible. The set on which the condition (7) is violated is a subset of the sector
@@
Let the expression $\|T^{t}\omega\|^{2}$ possess the angular distribution
$\mathbb{P}(S(\delta))=\int_{-\infty}^{\delta}f(t)\,\mathrm{d}t.$ (9)
Consider the solution $U$ for a rotationally symmetric right-hand side $F$ with finite norm $\|F\|_{s}$, and let $\chi(t)=1$ for $t\leq\delta$ and $\chi(t)=0$ for $t>\delta$. Then
$\int_{S(\delta)}|\widehat{U}(\omega)|\|\omega\|^{s+2}\,\mathrm{d}\omega\,\leq\frac{1}{n\nu_{n}}\int_{S^{n-1}}\frac{\chi\big{(}\|T^{t}\eta\|^{2}\big{)}}{\|T^{t}\eta\|^{2}}\,\mathrm{d}\eta\,\|F\|_{s}$ (10)
holds, independent of $\mu$ and with equality for $\mu=0$. This follows from the radial-angular representation of the involved integrals. By Lemma 2, the integral over the unit sphere on the right-hand side has a finite value for dimensions $m\geq 3$. In terms of the distribution density from (9), it is given by
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}\frac{\chi\big{(}\|T^{t}\eta\|^{2}\big{)}}{\|T^{t}\eta\|^{2}}\,\mathrm{d}\eta\,=\int_{0}^{\delta}\frac{f(t)}{t}\,\mathrm{d}t$ (11)
and tends to zero as $\delta$ goes to zero, for higher dimensions usually very rapidly as the example of the (rescaled) orthogonal projections from Theorem 4.1 shows.
The estimate (6) is extremely robust in many respects. First, it is based on a pointwise estimate of the Fourier transform of the error. It is therefore equally valid for other Fourier-based norms. Second, the function (3) can be replaced by any approximation $\widetilde{\alpha}(\omega)$ for which an estimate
@@
holds for the frequency vectors $\omega$ outside the sector (8). Therefore, the part of the solution associated with the $\omega\notin S(\delta)$ outside a sufficiently large ball around the origin can be neglected. In the high dimensions considered, the same holds for the part of the solution associated with the $\omega\notin S(\delta)$ in a sufficiently small ball around the origin, provided that the Fourier transform of $F$ can be controlled there.
Nevertheless, the estimate from Theorem 5.2 is rather pessimistic, because it only takes into account the decay of the left tail of the distribution (9), but ignores the fast decay of its right tail. To see what can be reached, we study the behavior of the iteration in the limit $\mu=0$, the case where the underlying effects are most clearly brought to light. Decomposing the vectors $\omega=r\eta$ into a radial part $r\geq 0$ and an angular part $\eta\in S^{n-1}$, the error (2) propagates frequency-wise as
@@
If the solution $U$ lies in the Barron space $\mathcal{B}_{s}(\mathbb{R}^{n})$, $s\geq 0$, this also holds for the iterates $U_{k}$ implicitly given by (16). The norm (15) of the iteration error then tends to zero for suitably chosen polynomials $P_{k}$ as $k$ goes to infinity.
@@
$\phi(\eta)\,=\,n\nu_{n}\int_{0}^{\infty}|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r.$
@@
Under a seemingly harmless additional assumption, one gets a rather sharp, hardly to improve estimate for the speed of convergence.
@@
Under the assumption that the norm given by (17) of the solution $U$ takes a finite value, the iteration error can be estimated as
@@
$\phi(\eta)\,=\,n\nu_{n}\int_{0}^{\infty}|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r$
@@
If one rewrites the integral still in terms of the density $f$, (17) follows. ∎
In sharp contrast to Theorem 5.1 and Theorem 5.2, this theorem strongly depends on the involved norms. But of course one can hope that other error norms behave similarly, especially for solutions whose Fourier transform is not strongly concentrated on a small sector around the kernel of $T^{t}$. For rotationally symmetric solutions $U$,
$\|U-U_{k}\|_{s}^{2}\leq\int_{-\infty}^{\infty}P_{k}(t)^{2}f(t)\,\mathrm{d}t\;\|U\|_{s}^{2}$ (18)
holds as the norms given by (15) and (17) coincide in this case. This estimate can in turn be transferred to the Hilbert spaces $H^{s}$, in which even equality holds.
@@
Let the rotationally symmetric function $U$ lie in the Hilbert-space $H^{s}$ equipped with the seminorm $|\cdot|_{2,s}$. The iteration error, here given by

### v5

@@ 4 The iterative procedure
and the functions $P_{k}(\lambda)$ are polynomials of order $k$ with value $P_{k}(0)=1$. Throughout this section, we assume that the expression $\|T^{t}\eta\|^{2}$ possesses the expected value one. We restrict ourselves to the analysis of this iteration in the spaces of the functions with finite norms (17). In parts, this analysis can be transferred to the Hilbert spaces $H^{s}$.
@@
If the solution $U$ possesses a finite norm $\|U\|_{s}$, this also holds for the iterates $U_{k}$. For all coefficients $\mu\geq 0$, the corresponding norm of the error, given by
$\|U-U_{k}\|_{s}=\int\|T^{t}\omega\|^{s}\,\big{|}P_{k}\big{(}\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)\big{)}\widehat{U}(\omega)\big{|}\,\mathrm{d}\omega,$ (4)
@@
be the to the interval $a\leq\lambda\leq b$ transformed Chebyshev polynomial $T_{k}$ of the first kind of degree $k$. The norm of the iteration error (2) then satisfies the estimate
$\|U-U_{k}\|_{s}\leq\frac{2q^{k}}{1+q^{2k}}\,\|U\|_{s}+\|U-U_{\delta}\|_{s},\quad q=\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1},$ (6)
@@
$\|T^{t}\omega\|^{2}>\delta\|\omega\|^{2}$ (7)
@@
It is $a\leq\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)\leq b$ if (7) holds and $0\leq\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)\leq b$ on the entire $\mathbb{R}^{n}$. Therefore, the proposition follows from the estimates
$|P_{k}(\lambda)|\leq 1\text{\;for $0\leq\lambda\leq a$},\quad|P_{k}(\lambda)|\leq\frac{2q^{k}}{1+q^{2k}}\text{\;for $a\leq\lambda\leq b$}$
@@
Depending on the size of $\kappa=\|T^{t}\|^{2}/\delta$, the norm of the error $U-U_{k}$ soon reaches the size of the norm of $U-U_{\delta}$. The idea behind the estimate (6) is that in high space dimensions the condition (7) is satisfied for nearly all $\omega$ and that the part $U-U_{\delta}$ of the solution $U$ is therefore negligible. Let $\phi:S^{n-1}\to\mathbb{R}$ be the integrable function
$\phi(\eta)=\,n\nu_{n}\int_{0}^{\infty}\|T^{t}\eta\|^{s}\,|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r$ (8)
from the definition (16) of the norms and let $S(\delta)$ be the sector
@@
With $M=S(\delta)\cap S^{n-1}$, then the remainder possesses the representation
$\|U-U_{\delta}\|_{s}=\frac{1}{n\nu_{n}}\int_{M}\phi(\eta)\,\mathrm{d}\eta.$ (10)
If $\phi$ is square integrable, that is, if the norm $|\!|\!|U|\!|\!|_{s}$ of the solution $U$ is finite, the remainder can be estimated by means of the Cauchy-Schwarz inequality. Expressing the area measure $\mathbb{P}(M)$ in terms of the density (9), this leads to
$\|U-U_{\delta}\|_{s}^{2}\leq\int_{0}^{\delta}f(t)\,\mathrm{d}t\;|\!|\!|U-U_{\delta}|\!|\!|^{2}_{s}.$ (11)
The right-hand side of this inequality tends at least like $\mathcal{O}(\delta^{m/2})$ to zero as $\delta$ goes to zero. The speed of convergence doubles if the function (8) remains bounded.
The estimate (6) is extremely robust in many respects. First, it is based on a pointwise estimate of the Fourier transform of the error. It is therefore equally valid for other Fourier-based norms. Second, the function (3) can be replaced by any approximation $\widetilde{\alpha}(\omega)$ for which, with an appropriate $\varepsilon>0$, an estimate
@@
holds for the frequency vectors $\omega$ outside the sector (9). Therefore, the part of the solution associated with the $\omega\notin S(\delta)$ outside a sufficiently large ball around the origin can be neglected. In the high dimensions considered, the same holds for the part of the solution associated with the $\omega$ in a small ball around the origin. The contribution of a ball $B(R)$ of radius $R$ around the origin to the norm (17) is
$\int_{B(R)}\|T^{t}\omega\|^{s}\,|\widehat{U}(\omega)|\,\mathrm{d}\omega\,=\,n\nu_{n}\int_{0}^{R}\psi(r)\,r^{s+n-1}\,\mathrm{d}r,$ (15)
where the function $\psi$ is the mean value
$\psi(r)=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\|T^{t}\eta\|^{s}\,|\widehat{U}(r\eta)|\,\mathrm{d}\eta.$ (16)
Provided the behavior of this mean value can be kept under control, the contribution of the ball rapidly decreases as soon as its radius falls below a certain bound.
Nevertheless, the estimate from Theorem 4.2 is rather pessimistic, because it only takes into account the decay of the left tail of the distribution, but ignores the fast decay of its right tail. To see what can be reached, we study the behavior of the iteration in the limit $\mu=0$, the case where the underlying effects are most clearly brought to light. Decomposing the vectors $\omega=r\eta$ into a radial part $r\geq 0$ and an angular part $\eta\in S^{n-1}$, the error (2) propagates frequency-wise as
@@
If the solution $U$ possesses a finite norm $\|U\|_{s}$, this also holds for the iterates $U_{k}$ implicitly given by (18). The norm (17) of the iteration error then tends to zero for suitably chosen polynomials $P_{k}$ as $k$ goes to infinity.
@@
$\phi(\eta)=\,n\nu_{n}\int_{0}^{\infty}\|T^{t}\eta\|^{s}\,|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r.$
@@
Under a seemingly harmless additional assumption, one obtains an estimate for the speed of convergence that can hardly be improved.
@@
Under the assumption that the norm $|\!|\!|U|\!|\!|_{s}$ of the solution $U$ takes a finite value, the iteration error can be estimated as
@@
$\phi(\eta)=\,n\nu_{n}\int_{0}^{\infty}\|T^{t}\eta\|^{s}\,|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r$
@@
If one rewrites the integral in terms of the density $f$, (19) follows. ∎
In contrast to Theorem 4.1 and Theorem 4.2, this theorem strongly depends on the involved norms. But of course one can hope that other error norms behave similarly.
@@
Let the function $U$ lie in the Hilbert-space $H^{s}$ equipped with the seminorm $|\cdot|_{2,\,s}$. The iteration error, for rotationally symmetric functions $U$ given by